OpenStudy (anonymous):

Use the distance formula to find the equation of a parabola with focus (-3, 0) and directrix x = 6. Thanks!

OpenStudy (anonymous):

I got lost at (

OpenStudy (anonymous):

by definition of a parabola, it is the locus of points that are equidistant from the focus and the directrix. let (x,y) be a point on the parabola 1) its distance from (-3,0) \(=\sqrt{[x-(-3)]^2+(y-0)^2}\) 2) its distance from the directrix x=6 \(\Large=\frac{|x+6|}{1}=|x+6|\) by definition, they are equal!

OpenStudy (anonymous):

\[ |x+6|=\sqrt{(x+3)^2+y^2} \] square both sides and simplify

OpenStudy (anonymous):

correction... it should be : \[|x-6|=\sqrt{(x+3)^2+y^2}\]

OpenStudy (anonymous):

and evaluate

OpenStudy (anonymous):

Okay... so, For the first part, I got -36, I don't think I understand how to do everthing under the square root sign.

OpenStudy (anonymous):

@electrokid

OpenStudy (anonymous):

just the 36? you are supposed to get an equation. it should look like x=..... something in "y" and a constant term

OpenStudy (anonymous):

Yeah, I did the absolute value stuff first @electrokid

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

you can take a picture of your work and post it here so I see

OpenStudy (anonymous):

-6 square rooted... WAIT wouldn't it be 36 not -36?

OpenStudy (anonymous):

I do not follow your statement. use the equation editor below to write what you did

OpenStudy (anonymous):

how did you get rid of the "absolute value" sign?

OpenStudy (anonymous):

\[|x-6| =x-6 which is 36\]

OpenStudy (anonymous):

but we do not have the "x"

OpenStudy (anonymous):

what formula or property did you use ?

OpenStudy (anonymous):

Nope its -36. you cant get rid of the sigh if its negative

OpenStudy (anonymous):

the only way to get rid of an absolute value sign is 1) by squaring it OR 2) by writing two possible results

OpenStudy (anonymous):

since we want only one equation, we square it so, we expand \[(x-6)^2\]

OpenStudy (anonymous):

Oh okay

OpenStudy (anonymous):

what do we get for the expansion?

OpenStudy (anonymous):